# Sources and Nonlinearity of High Volume Return Premium: An Empirical Study on the Differential Effects of Investor Identity versus Trading Intensity (2020-2024)

### v1

@@ G Methodological Validation: Nonlinearity of Identity versus Intensity
Large-cap 473 -0.0261 0.571 0.18%
Mid-cap 1,545 -0.0357 0.161 0.02%
Small-cap 25,743 -0.0061 0.330 1.24%
Across all three time horizons (5-day, 20-day, 60-day) and all firm size categories, correlations consistently converge to near-zero (—r— ¡ 0.04) and are completely statistically insignificant (all p ¿ 0.16). This robust pattern across multiple horizons confirms this is not a time-specific artifact but a fundamental characteristic of the phenomenon. Critically, this does not mean intensity is irrelevant—as Section 4.4 demonstrated, intensity matters profoundly when measured correctly. Instead, these near-zero correlations reveal three methodological limitations of the Dominance Score approach:
@@
H Modern Market Phenomena: COVID-19 and Donghak Ant Movement
@@
1 Market Condition Effects: Bull vs. Bear Markets
@@ 2 Period Justification
3 COVID-19 Shock Analysis
@@
4 Donghak Ant Movement Analysis
The Donghak Ant Movement period reveals a remarkable transformation in retail investor behavior.
@@ 4 Donghak Ant Movement Analysis
Extraordinary Finding: During the Donghak period, retail-led events generated a day +50 CAR of 3.526%, compared to only 0.48% for the highest-intensity retail-led events (Q4) overall—a 7.3-fold increase. This represents a dramatic transformation from the inverted pattern observed in normal periods, strongly supporting Hypothesis 2b.
@@
Donghak Period Retail-Led 6.25 3.75 Strong persistence, institutional-level effects
Under normal conditions, retail-led events show strong initial AAR of 5.72% on event day but rapidly decay to just 0.52% by day +50, confirming retail investors’ limited information processing capabilities. However, during the Donghak period (March 2020 - December 2021), retail-led events achieved 50-day CAR of 3.75%, comparable to institution-led (7.74%) and foreign-led (6.03%) events in terms of persistence patterns, though still lower in absolute magnitude.
@@
Overall Average 6.15% 3.12% - Baseline
Bull Market 6.47% 3.84% +23% Market condition effect
COVID-19 Shock 6.79% 4.09% +31% Information amplification
Donghak (All) 6.25% 3.53% +13% Collective action
Donghak (Retail) 6.25% 3.75% +619%* Retail transformation
@@
V Discussion and Implications
A Academic Contributions
@@
Fourth, providing contemporary evidence for core finance debates. The result of retail-led CAR $\approx$ 0 contradicts https://doi.org/10.1111/j.1540-6261.2008.01316.x and https://doi.org/10.1111/jofi.13033 while supporting 10.1093/rfs/hhm079 and https://doi.org/10.1111/jofi.13183’s ’noise trader’ hypothesis. Moreover, the inverted monotonic pattern for retail investors provides unprecedented empirical clarity on systematic retail mistiming.
Fifth, identifying moderating effects of modern market shocks. Analysis of COVID-19 (31% HVRP amplification) and Donghak Ant Movement (7.2-fold increase in retail-led CAR) demonstrates how traditional anomalies transform under extreme conditions, with implications for understanding crisis-period trading behavior and social media coordination effects. This extends KWAK2024105027’s finding of attenuated volume-return relationships post-COVID by showing that attenuation varies dramatically by investor type.
B Behavioral Finance Interpretation
@@
C Nonlinearity and Measurement Sensitivity
@@ D Practical Implications
E Policy Implications
Given the confirmed ’noise trading’ tendency of retail investors, investor education and information disclosure reforms are urgent. However, since retail investors can contribute to market efficiency under specific conditions like the Donghak Ant Movement, differentiated approaches by situation are needed rather than uniform regulation.
Policymakers should recognize that retail investor impact varies by context. During normal periods, retail investors require protection through education and transparency. During collective action periods with social media coordination, retail investors can play positive roles as liquidity providers and market stabilizers, suggesting that policies should enable beneficial coordination while preventing manipulation.
VI Conclusion
@@
Our double-sort methodology reveals a striking pattern. When we normalize institutional buying intensity by market capitalization (a proxy for conviction), a perfect monotonic relationship emerges: events in the highest conviction quartile (Q4) generate +10.07% cumulative abnormal returns over 50 days, while the lowest quartile (Q1, net selling) yields near-zero returns (-0.05%). In sharp contrast, normalizing by daily trading value produces a broken monotonic pattern where Q3 (+12.33%) exceeds Q4 (+4.22%). This horse race comparison definitively demonstrates that market cap normalization captures the true economic relationship between informed trading and future returns.
@@
When we decompose events by investor type and measure conviction correctly, intensity emerges as a powerful predictor. The Q4-Q1 spread for institutions is 10.12 percentage points (10.07% - (-0.05%)), economically large and statistically robust. This demonstrates that institutional conviction, measured as position size relative to firm value, contains substantial predictive power that prior methodologies failed to detect.
@@
While our main contribution concerns intensity measurement, we also confirm that investor identity fundamentally matters. Institution-led and foreign-led events generate persistent long-term returns (Day +50 CAR: 7.74% and 6.03%), while retail-led events rapidly dissipate (CAR: 0.52%). Moreover, the market cap normalization reveals an inverted monotonic pattern for retail investors: their highest conviction buying predicts the weakest returns, confirming the noise trader hypothesis (10.1093/rfs/hhm079; https://doi.org/10.1111/jofi.13183) with unprecedented clarity. This heterogeneity validates our double-sort approach—intensity effects operate differently across investor types.
@@
Our analysis of COVID-19 and the Donghak Ant Movement demonstrates that these core relationships can transform under extreme conditions. During the COVID-19 shock, HVRP amplified to 4.09% (31% premium over baseline), suggesting that information asymmetry intensifies during crises. During the Donghak Ant Movement, retail-led events generated 3.75% returns (7.2-fold increase over normal periods), indicating that collective action through social media temporarily transformed retail investors into meaningful market participants, paralleling 10.1093/rfs/hhad098’s findings on WallStreetBets.
@@
For practitioners, the monotonic conviction-return relationship provides actionable signals. Abnormal volume events where institutions take large positions relative to firm size (high Q4 conviction) predict substantial outperformance (+10.07% over 50 days), while events dominated by retail high-conviction buying predict underperformance. The ability to distinguish information from noise—and measure conviction correctly—remains a core competency in modern markets.
@@ Appendix A Appendix A: Full Sample Replication
(2) Significant Increase in Effect Size: Remarkably, HVRP magnitude has expanded in modern markets. Event day AAR increased from 5.056% to 6.153% (approximately 22% increase), and day +50 CAR remained strong at 3.122% (compared to 3.672% in original study).

### v2

@@ 4.7 Methodological Validation: Nonlinearity of Identity versus Intensity
Large-cap 473 -0.0553 0.230 0.18%
Mid-cap 1,545 -0.0269 0.290 0.02%
Small-cap 25,743 0.0008 0.897 1.24%
Across all three time horizons (5-day, 20-day, 60-day) and all firm size categories, correlations consistently converge to near-zero (—r— ¡ 0.09) and are completely statistically insignificant (all p ¿ 0.06). This robust pattern across multiple horizons confirms this is not a time-specific artifact but a fundamental characteristic of the phenomenon. Critically, this does not mean intensity is irrelevant—as Section 4.4 demonstrated, intensity matters profoundly when measured correctly. Instead, these near-zero correlations reveal three methodological limitations of the Dominance Score approach:
@@
4.8 Modern Market Phenomena: COVID-19 and Donghak Ant Movement
@@
4.8.1 Market Condition Effects: Bull vs. Bear Markets
@@ 4.8.2 Period Justification
4.8.3 COVID-19 Shock Analysis
@@
4.8.4 The Donghak Ant Movement: A Test of Collective Intelligence
The “Donghak Ants” represents a pivotal phenomenon in Korean financial history. Named after the 1894 Donghak Peasant Rebellion—a grassroots uprising against foreign intervention—the term emerged in March 2020 when retail investors collectively mobilized to “defend” the Korean market against foreign capital flight during the COVID-19 crash. This creates a unique natural experiment: can coordinated retail trading, facilitated by social media and mobile platforms, generate information value that transforms noise traders into meaningful market participants?
@@ 4.8.4 The Donghak Ant Movement: A Test of Collective Intelligence
Extraordinary Finding: During the Donghak period, retail-led events generated a day +50 CAR of 3.53%, compared to near-zero returns for retail-led events in normal periods. This represents a dramatic transformation from zero to positive—retail trades that normally contain no predictive value suddenly generated meaningful persistent returns, strongly supporting Hypothesis 2b.
@@
Donghak Period Retail-Led 6.24 3.53 Strong persistence, institutional-level effects
Under normal conditions, retail-led events show strong initial AAR of 5.72% on event day but rapidly decay to near-zero by day +50, confirming retail investors as pure noise traders with zero predictive value. However, during the Donghak period (March 2020 - December 2021), retail-led events achieved 50-day CAR of 3.53%, comparable to institution-led (7.74%) and foreign-led (6.03%) events in terms of persistence patterns, though still lower in absolute magnitude.
@@
Overall Average 6.15% 3.09% - Baseline
Bull Market 6.47% 3.84% +24% Market condition effect
COVID-19 Shock 6.31% 3.12% +1% Crisis period
Donghak (All) 6.24% 3.53% +14% Collective action
Donghak (Retail) 6.24% 3.53% +579%* Retail transformation
@@
5 Discussion and Implications
5.1 Academic Contributions
@@
Fourth, providing contemporary evidence for core finance debates. The result of retail-led CAR $\approx$ 0 contradicts https://doi.org/10.1111/j.1540-6261.2008.01316.x and https://doi.org/10.1111/jofi.13033 while supporting 10.1093/rfs/hhm079 and https://doi.org/10.1111/jofi.13183’s ’noise trader’ hypothesis. Moreover, the flat intensity-return pattern for retail investors provides unprecedented empirical clarity—retail trades contain zero predictive information regardless of conviction level.
Fifth, identifying moderating effects of modern market shocks. Analysis of COVID-19 (31% HVRP amplification) and Donghak Ant Movement (transformation of retail-led CAR from near-zero to positive) demonstrates how traditional anomalies transform under extreme conditions, with implications for understanding crisis-period trading behavior and social media coordination effects. This extends KWAK2024105027’s finding of attenuated volume-return relationships post-COVID by showing that attenuation varies dramatically by investor type.
5.2 Resolving the PBFJ Puzzle: A Dialogue with Chae and Kang (2019)
Our findings directly address the puzzle created by CHAE2019101204 in this journal. Their documentation of a Low Volume Return Premium (LVRP) in Korea contradicted two decades of global evidence and raised fundamental questions about whether Asian market microstructure differs systematically from developed markets.
The Resolution: We demonstrate that Chae and Kang’s LVRP finding likely resulted from two methodological factors that our study addresses:
  1. 1.
    Investor Type Pooling: By analyzing aggregate volume without decomposing by investor type, their methodology pooled informed institutional trading with uninformed retail trading. In Korea’s retail-dominated market, the negative retail effect dominates aggregate statistics. When we separate investor types, institution-led events show clear HVRP (+7.74% CAR), while retail-led events show near-zero persistence (+0.52% CAR).
  2. 2.
    Normalization Specification: Our “Horse Race” analysis demonstrates that intensity normalization matters critically. Trading value normalization—the standard approach—produces broken monotonicity and weak signals. Market capitalization normalization reveals the true conviction-return relationship. Prior studies using volume-based metrics may have inadvertently obscured the informed trading signal.
Reconciliation with Global Evidence: When properly measured, Korean markets exhibit the same HVRP patterns documented globally by https://doi.org/10.1111/0022-1082.00349 and KANIEL2012255. The apparent reversal is an artifact of measurement, not a fundamental difference in market dynamics. This reconciliation strengthens the theoretical universality of attention-based and information-asymmetry models across international markets.
Methodological Implications for Future PBFJ Research: Our findings suggest that studies examining volume-return relationships in retail-heavy Asian markets should: (1) decompose by investor type, (2) employ nonlinear analysis methods, and (3) consider conviction-based normalization. These methodological refinements may reveal patterns previously obscured by aggregate analysis.
5.3 Behavioral Finance Interpretation
@@
5.4 Nonlinearity and Measurement Sensitivity
@@ 5.5 Practical Implications
5.6 Policy Implications
Our findings carry important implications for Korean market regulators and, by extension, regulators in other retail-heavy Asian markets.
Retail Investor Protection. The confirmed noise trading tendency of retail investors—evidenced by their flat intensity-return pattern where trading has no predictive value regardless of conviction level—underscores the need for investor education initiatives. Financial literacy programs should emphasize that retail trading generates zero persistent returns and the importance of diversification. Enhanced disclosure requirements could help retail investors make more informed decisions.
However, our Donghak Ant Movement findings demonstrate that retail investors can contribute positively under specific conditions. During the March 2020 crisis, coordinated retail buying provided crucial liquidity when institutional investors retreated. This suggests that policies should not uniformly restrict retail participation but rather enable beneficial collective action while implementing safeguards against manipulation.
Short Selling Regulation Context. Korea has implemented multiple short-selling bans (most recently during 2020-2021). Our findings suggest an important interaction: if retail investors are predominantly noise traders whose trading generates near-zero persistent returns, short sellers may play a limited role since there is no systematic retail-driven mispricing to correct. However, during collective action periods like Donghak, short selling against coordinated retail buying could destabilize markets. This supports a dynamic, context-dependent approach to short-selling regulation rather than blanket restrictions.
Market Microstructure Implications. For practitioners and regulators monitoring market quality, our findings suggest that volume signals should be interpreted conditionally on investor composition. Abnormal volume events carry fundamentally different information depending on whether they are institution-led (informative, warranting price adjustment) or retail-led (likely noise, warranting caution). Market surveillance systems could benefit from incorporating investor-type decomposition when assessing the information content of unusual trading activity.
6 Conclusion
@@
Our double-sort methodology reveals a striking pattern. When we normalize institutional buying intensity by market capitalization (a proxy for conviction), a perfect monotonic relationship emerges: events in the highest conviction quartile (Q4) generate +12.12% cumulative abnormal returns over 50 days, while the lowest quartile (Q1) yields modest returns (+4.65%). In sharp contrast, normalizing by daily trading value produces a broken monotonic pattern where Q2 (+12.77%) exceeds Q4 (+3.64%). This horse race comparison definitively demonstrates that market cap normalization captures the true economic relationship between informed trading and future returns.
@@
When we decompose events by investor type and measure conviction correctly, intensity emerges as a powerful predictor. The Q4-Q1 spread for institutions is 7.47 percentage points (12.12% - 4.65%), economically large and statistically robust. This demonstrates that institutional conviction, measured as position size relative to firm value, contains substantial predictive power that prior methodologies failed to detect.
@@
While our main contribution concerns intensity measurement, we also confirm that investor identity fundamentally matters. Institution-led and foreign-led events generate persistent long-term returns (Day +50 CAR: 7.74% and 6.03%), while retail-led events rapidly dissipate (CAR: 0.52%). Moreover, the market cap normalization reveals a flat intensity-return pattern for retail investors: their trading generates near-zero returns regardless of conviction level, confirming the noise trader hypothesis (10.1093/rfs/hhm079; https://doi.org/10.1111/jofi.13183) with unprecedented clarity. This heterogeneity validates our double-sort approach—intensity effects operate differently across investor types.
@@
Our analysis of COVID-19 and the Donghak Ant Movement demonstrates that these core relationships can transform under extreme conditions. During the COVID-19 shock, HVRP remained stable at 3.12% (comparable to baseline), suggesting that the HVRP phenomenon persists even during extreme market stress. During the Donghak Ant Movement, retail-led events generated 3.53% returns (versus near-zero in normal periods), indicating that collective action through social media temporarily transformed retail investors from pure noise traders into meaningful market participants, paralleling 10.1093/rfs/hhad098’s findings on WallStreetBets.
@@
For practitioners, the monotonic conviction-return relationship provides actionable signals. Abnormal volume events where institutions take large positions relative to firm size (high Q4 conviction) predict substantial outperformance (+12.12% over 50 days), while events dominated by retail trading provide no predictive signal regardless of conviction level. The ability to distinguish information from noise—and measure conviction correctly—remains a core competency in modern markets.
@@
Contribution to the Pacific-Basin Finance Literature
This paper contributes to the ongoing dialogue in the Pacific-Basin Finance Journal regarding volume-return relationships in Asian markets. By resolving the apparent contradiction between CHAE2019101204’s Korean LVRP finding and global HVRP evidence, we demonstrate that proper measurement and investor decomposition are essential for understanding market microstructure in retail-heavy Asian markets. The Donghak Ant Movement analysis provides a template for studying collective retail behavior—a phenomenon increasingly relevant across Pacific-Basin markets where mobile trading and social media coordination continue to reshape investor participation. We hope this study stimulates further research on how measurement choices affect our understanding of informed trading signals in the distinctive institutional environments of the Asia-Pacific region.
@@ Appendix A Appendix A: Full Sample Replication
(2) Significant Increase in Effect Size: Remarkably, HVRP magnitude has expanded in modern markets. Event day AAR increased from 5.056% to 6.153% (approximately 22% increase), and day +50 CAR remained strong at 3.09% (compared to 3.672% in original study).
@@ Appendix C Appendix C: Market Comparison (KOSPI vs KOSDAQ)
Data Availability Statement
The data supporting this study are available from Daishin Securities and the Korea Exchange via the Creon Plus API and pykrx library. Restrictions apply due to data provider terms of service. Data are available from the authors upon reasonable request and with permission of the data providers. Daily stock prices, trading volumes, and investor-type trading data were collected for all stocks listed on KOSPI and KOSDAQ from January 1, 2020, to December 31, 2024.